# Spontaneous phase locking in a broad-area semiconductor laser

### v1

@@ III Discussion
Our experiments reveal a spontaneous phase locking of transverse modes in a simple broad-area Fabry-Perot cavity without special features or external control elements. This is surprising given the significant spatio-temporal instabilities of such devices. Since only two out of up to eight transverse mode orders are phase locked, the dynamics of the BAL remains complex and unstable. Hence the phase-locking effect is not easily recognized, and direct measurements of the phase fluctuation correlations are needed to ascertain which modes are locked.
@@
Mode locking can be induced by various techniques including saturable absorbers, external modulation and feedback. Longitudinal [33] and transverse mode-locking [34] in broad-area lasers has also been induced via pulsed injection. Furthermore, spatio-temporal mode locking has been demonstrated for Helium-Neon lasers [46], Ti:Sapphire lasers [47] and fiber lasers [15, 16], but requires careful design of the laser system and its cavity. In our case, in contrast, phase locked multiplets emerge spontaneously from the internal laser dynamics of a BAL with a normal Fabry-Perot cavity. While spontaneous phase, frequency and mode locking have been observed before for semiconductor lasers, our case is different to previous works in several respects as discussed in the following.
@@
For semiconductor lasers with relatively narrow FP cavities featuring only the 1st- and 2nd-order transverse modes, phase locking between the transverse modes has been observed [48], which leads to a periodic oscillation of the beam profile, as well as frequency locking [49, 50], which results in asymmetric, static emission profiles and beam steering. Phase locking [51, 52] and frequency locking [53, 54] between two different spatial modes have also been observed for asymmetric cavities. However, in these examples only two different spatial modes lase in the first place, whereas our BAL features up to $8$ transverse lasing modes. Moreover, only modes that naturally exist in the passive cavity were locked, whereas we observe multiplets of modes newly created by the nonlinear dynamics.
@@
Quantum cascade lasers (QCLs) can exhibit spontaneous mode locking between longitudinal modes [55, 56]. Furthermore, QCL frequency combs with phase and frequency locking of several transverse modes have been demonstrated [57, 58]. However, the dynamics of QCLs is quite different from that of quantum well lasers due to the much shorter gain recovery time and the strong third-order nonlinearity [55] of the former. Locking processes in QCLs are mainly attributed to four-wave mixing, whereas, as we concluded earlier, the creation and phase locking of transverse modes presented here cannot be explained by four-wave mixing.
The most interesting aspect of our laser is the coexistence of groups of coherent modes (the phase locked multiplets) with a large number of incoherent (not phase locked) modes, reminiscent of chimera states [21, 22]. Chimera states have been reported for semiconductor lasers with feedback [24, 25], in which case virtual oscillators in the feedback loop are considered instead of actual lasing modes like in our case. Spontaneous synchronization of mode clusters was found for a quantum dot laser [27], but the synchronization concerned the amplitude dynamics of different longitudinal modes, not the phases. Cluster synchronization was furthermore found among the longitudinal modes of a semiconductor laser with absorbing section [26], and among the beat notes of a QCL with tunable mirror reflectivity [28]. While Refs. [27, 26, 28] show synchronization between different longitudinal modes, which can be considered as identical oscillators, we demonstrate for the first time spontaneously-formed mode clusters with different transverse modes, that is, clusters of non-identical oscillators. More importantly, in our case the nonlinear dynamics is not only responsible for coupling different oscillators as in previous works, but it moreover creates additional oscillators that are not present in the system a priori.
@@
The observation of groups of phase locked modes in a commercial semiconductor laser indicates that cluster synchronization and chimera states may be more frequent in lasers and other nonlinear systems than previously thought. The spatio-temporal dynamics of broad-area lasers is still not fully understood, and our observation of spontaneous phase locking gives a new perspective for understanding and possibly controlling their instabilities using, e.g., optical injection or special cavity designs. This will have important implications for our fundamental understanding of complex spatio-temporal and multimode dynamics as well as for applications like mode-locked lasers.
@@ Acknowledgements.
The Chair in Photonics is supported by Région Grand Est, GDI Simulation, Département de la Moselle, European Regional Development Fund, CentralSupélec, Fondation CentralSupélec, and the Eurometropole de Metz. S.B. thanks Mario Fernandes for technical support.
@@ A.1 Broad-area laser
We use a commercial quantum-well semiconductor laser from Sheaumann (P/N CM-850-1000-050) mounted on a C-mount. The cavity is laterally defined by etching to just above the active layer and depositing a dielectric afterwards, which gives the laser an aspect of both gain and index guiding. The laser is operated in a temperature-controlled laser diode mount (Thorlabs LDMC20/M) and pumped in continuous wave operation by a diode driver (Thorlabs ITC4005). The temperature of the mount is kept constant at $21^{\circ}$ during all measurements unless otherwise noted. The laser emission is transverse-electric polarized (electric field parallel to the cavity plane).
@@
A.2 Time-domain and spectral measurements
Time-domain measurements were performed with a fast photodetector (Newport 1484-A-50, $15$ kHz to $22$ GHz) connected to a traveling-wave amplifier (Newport 1421, $8.5$ dB) and an oscilloscope (Tektronix DPO72304SX, $23$ GHz). Time traces were measured with $50$ GSample/s ($20$ ps time step) and $10^{6}$ points ($20~{}\mu$s long). RF-spectra were obtained by calculating the Fourier transform of the time traces and were smoothed with a $2$ MHz wide moving average. It should be noted that extremely narrow peaks are observed in many RF-spectra: these are artifacts due to synchronization issues between different analog-to-digital converters of the oscilloscope and appear even without input signal. Hence they have no physical significance and should be ignored.
For measurements of the total laser emission, the beam was coupled into a graded-index multimode fiber (MMF, $50~{}\mu$m diameter, NA$=0.20$, Thorlabs M116L02) using a 20×20$\times$20 × aspheric objective lens (Newport 5724-B-H, NA = $0.5$) in a fiber coupler. Feedback from reflections at the fiber facet were avoided via an optical isolator (OI, Thorlabs IO-5-850-VLP). The multimode fiber was connected to the photodetector or the OSA. It should be noted that coupling the beam of a multi-transverse mode laser into a single-mode fiber is very inefficient, and more importantly the higher-order transverse modes have a lower coupling efficiency compared to low-order modes, which results in an observational bias.
Spatio-spectral images with low resolution (Fig. 1c) were obtained with an imaging spectrometer (Princeton Instruments SpectraPro HRS-500 with $1800$ g/mm grating, $30$ pm resolution) and a CCD camera (Allied Vision Mako U-503B). In order to measure the spectrally-resolved near-field intensity distributions, the output facet of the BAL was imaged onto the entrance slit of the spectrometer with 40×40$\times$40 × magnification. A dove prism (Thorlabs PS992M-B) was used to rotate the image of the BAL facet by $90^{\circ}$ so it matches the vertical orientation of the entrance slit.
A.3 Low-resolution optical spectra
Figure 5 shows low-resolution lasing spectra of the BAL measured with an optical spectrum analyzer (OSA, Anritsu MS9740A) with $30$ pm resolution ($12.5$ GHz). While the spectral resolution of the OSA is relatively low, it allows measurements over a broader spectral range than the heterodyning technique. At low resolution, the spectra feature a single series of peaks with spacing equal to the free spectral range (FSR) of $25.9$ GHz ($62.4$ pm), hiding the fact that the laser exhibits several transverse modes. As the pump current increases, the center of mass (COM) of the spectrum red-shifts (see Fig. 5b) due to Joule heating. The total shift is about $2$ nm over $800$ mA. The red-shift of the individual peaks in the spectrum can also be seen in 5a.
@@ A.3 Low-resolution optical spectra
A.4 Near- and far-field intensity distributions
The output facet of the BAL is imaged onto a CMOS camera (Allied Vision Mako U-503B) using a 25×25$\times$25 × microscope objective (NA $=0.5$) and a lens with $175$ mm focal length (Thorlabs LA1399-B) in a telescope configuration. A beam splitter and third lens (Thorlabs LA1050-B, $100$ mm focal length) are used to obtain the Fourier transform of the image plane for parallel far-field measurements with a second camera (Allied Vision Mako U-130B). The far-field angle is calculated from the magnification of the telescope and the focal length of the third lens. Neutral-density filters in the beam path and the exposure times of the cameras are adjusted during the measurement to avoid both low signal strength and saturation.
Figure 6a shows the measured near-field intensity distributions, which are almost uniformly flat well below threshold ($I_{th}=160$ mA). Just above threshold, the intensity distribution exhibits a strong dip in the center with two peaks near the edges of the output facet. As the pump current increases, these two peaks move further towards the edges, and the center of the intensity distribution slowly fills. Above approximately $850$ mA the near-field distribution develops a top hat profile with significant modulation (see Fig. 1b). A similar evolution of the near-field intensity distributions as function of the pump current was observed in Ref. [59] and attributed to the interplay of index guiding and carrier-induced index changes. It should be noted that the near-field images are very symmetric around $x=0$ even though the individual transverse mode profiles show significant asymmetries.
The far-field intensity distributions shown in Fig. 6b are very broad below threshold and collapse to a single lobe at $0^{\circ}$ just above threshold. In this current regime, mainly the 1st- and 2nd-order modes lase, but the near-field distributions exhibit a pronounced dip in the center nonetheless since the fundamental mode features a trough like the $q=2$ mode (see Fig. 2c). As the pump current increases, higher-order modes start to lase, and the power shifts away from the low order modes.
@@ B.1 Experimental setup and measurement protocol
Figure 7: Experimetnal setup for spatially-resolved heterodyne measurements. a, Free-space part of the setup, where $x$ indicates the lateral position of the fiber. BAL: broad-area laser, Obj: 40×40$\times$40 × objective lens, OI: optical isolator, MMF: multimode fiber. b, Fiber-optic part of the setup. Single-mode fibers (multimode fibers) are indicated in yellow (orange), and coaxial cables in blue. TL: tunable laser (reference laser), eVOA: electronic variable optical attenuator, $\lambda$-meter: wavelength meter, PD: fast photodetector, RF-amp: RF-amplifier, OSA: optical spectrum analyzer, scope: oscilloscope
Figure 7 shows the experimental setup for spatially-resolved heterodyne measurements. A 40×40$\times$40 × aspherical objective lens (Newport 5722-B-H, NA = $0.6$) and a plano-convex lens (Thorlabs LA1979-B, focal length $200$ mm) image the output facet of the BAL with 40×40$\times$40 × magnification in the plane where its emission is collected with a MMF (Thorlabs M116L02, see Appendix A.2). An optical isolator (Thorlabs IO-5-850-VLP) avoids parasitic feedback into the BAL from reflections at the fiber facet. The fiber is mounted on a motorized translation stage (Thorlabs NRT100/M with driver BSC201) and scans the transverse position on the output facet of the BAL with a step size of $20~{}\mu$m (141 positions). The corresponding step size on the BAL output facet, $0.5~{}\mu$m, is slightly smaller than the spatial resolution of $0.7~{}\mu$m estimated from the NA of the objective lens.
The signals from the BAL and the reference laser (tunable laser Toptica DL pro with Diode AR 860 nm) are superimposed via a multimode fiber-optic coupler (Thorlabs TM50R5F2B) on a fast photodetector and measured with an oscilloscope (see Appendix A.2). The optical powers of the reference laser and the BAL on the photodetector are adjusted to about $0.75$ mW and $0.5$ mW, respectively, using a neutral density filter and a variable optical attenuator (VOA), though the power of the BAL can be smaller for low pump currents. The reference laser is set to $20^{\circ}$ with $143$ mA pump current and its wavelength adjusted to the desired value. It is left to stabilize for at least $20$ min, and we check that it is in a stable single-mode lasing regime. Its wavelength is monitored using a wavelength meter (High Finesse Angstrom WS6-600 IRVIS) connected via a $99:1$ fiber coupler (Thorlabs TW850R1A2).

### v2

@@ III Discussion
Our experiments reveal a spontaneous phase locking of transverse modes in a simple broad-area Fabry-Perot laser without special features or external control elements. This is surprising given the significant spatio-temporal instabilities of such devices. Since only two out of up to eight transverse mode orders exhibit phase locking, the dynamics of the BAL remains complex and unstable. Hence the phase-locking effect is not easily recognized, and direct measurements of the phase fluctuation correlations are needed to ascertain which modes are locked.
@@
Mode locking can be generated by techniques like saturable absorbers, external modulation and feedback, and pulsed injection into broad-area lasers can induce longitudinal [35] and transverse mode-locking [36]. Furthermore, spatio-temporal mode locking has been demonstrated for Helium-Neon lasers [47], Ti:Sapphire lasers [48] and fiber lasers [15, 16], but requires careful design of the laser system and its cavity. Here, in contrast, phase-locked multiplets emerge spontaneously from the intrinsic laser dynamics of a BAL with a normal Fabry-Perot cavity. While spontaneous phase, frequency and mode locking have been previously observed for semiconductor lasers, our case is different in several respects.
@@
For semiconductor lasers with quite narrow FP cavities featuring only the 1st- and 2nd-order transverse modes, phase locking between the transverse modes has been observed [49], which leads to a periodic oscillation of the beam profile, as well as frequency locking [50, 51], which results in asymmetric, static emission profiles. Phase locking [52, 53] and frequency locking [54, 55] between two different spatial modes have also been observed for asymmetric cavities. However, in these examples only two different spatial modes lase in the first place, whereas our BAL features up to eight transverse lasing modes. Moreover, previously only modes that naturally exist in the passive cavity were locked, whereas we observe multiplets of modes created by the nonlinear dynamics.
@@
Spontaneous mode locking between longitudinal modes has been observed for quantum-dot [56], quantum-dash [57] and quantum-cascade lasers (QCLs) [58, 59]. Furthermore, QCL frequency combs with phase and frequency locking of several transverse modes have been demonstrated [60, 61]. However, the dynamics in these lasers is quite different from that of quantum well lasers due to the much shorter gain recovery time [62] and the strong third-order nonlinearity of the former [58]. Locking of transverse modes in QCLs is primarily attributed to four-wave mixing [60, 61], whereas our observations cannot be explained by four-wave mixing since the multiplet modes are not equidistant.
Figure 5: Global degree of synchronization. a, Total number of measured modes (blue), number of modes with transverse quantum numbers $q=1$ and $2$ (red), and number of phase-locked modes in multiplets (yellow). b, Fraction of modes with transverse quantum numbers $q=1$ and $2$ (red) and of phase-locked modes in multiplets (yellow). The fraction of phase-locked modes can be considered as an order parameter indicating the degree of global synchronization. In contrast, the LRF in Fig. 3e represents the degree of local synchronization, which shows a qualitatively different dependency on the pump current.
[…]
Chimera states have been reported for semiconductor lasers with feedback [25, 26], where virtual oscillators in the feedback loop are considered. Cluster synchronization were furthermore found in quantum-dot [27] and quantum-cascade [28] lasers with absorbing sections. Our system is different in several respects. Partial synchronization is not induced by feedback or absorbing sections, but results from the inherent laser dynamics, and is formed between different transverse modes. In contrast to Refs. [25, 26], we observe synchronization between actual lasing modes instead of virtual oscillators. Moreover, some of the synchronized lasing modes do not exist in the system a priori, but are created by the nonlinear dynamics, another distinction from previous works.
@@
Our observation of groups of phase-locked modes in a commercial semiconductor laser indicates that partially synchronized dynamics may be more frequent in lasers and other nonlinear systems than previously thought. The spatio-temporal dynamics of broad-area lasers is still not fully understood, and our observation of spontaneous phase locking gives a new perspective for understanding and possibly controlling their instabilities using, e.g., optical injection or special cavity designs. Figure 5 shows that the pump current determines the degree of synchronization, but further studies are needed to understand the underlying physical mechanisms. This will have important implications for our fundamental understanding of complex spatio-temporal and multimode dynamics as well as for applications like spatio-temporal mode locking and high-power lasers.
@@ Acknowledgements.
The Chair in Photonics is supported by Région Grand Est, GDI Simulation, Département de la Moselle, European Regional Development Fund, CentraleSupélec, Fondation CentraleSupélec, and the Eurometropole de Metz. S.B. thanks Mario Fernandes for technical support.
@@ A.1 Broad-area laser
We use a commercial quantum-well semiconductor laser (Sheaumann Laser, P/N CM-850-1000-050) soldered to a C-mount. The cavity is laterally defined by etching to just above the active layer and depositing a dielectric afterwards, which gives the laser an aspect of both gain and index guiding. The laser is operated in a temperature-controlled laser diode mount (Thorlabs LDMC20/M) and pumped in continuous wave operation by a diode driver (Thorlabs ITC4005). The temperature of the mount is kept constant at $21^{\circ}$ during all measurements unless otherwise noted. The laser emission is transverse-electric polarized (electric field parallel to the cavity plane).
Figure 6: LI-curve of the broad-area laser. The LI-curve was measured in free space and shows a threshold at $160$ mA.
Figure 6 shows the LI-curve of the broad-area laser. It was measured in free space by using an objective and a lens to collect and focus the laser emission on a power meter (Newport 818-IG). The threshold is at $I_{th}=160$ mA. The maximal measured power is in good agreement with the specifications ($1$ W at $1100$ mA).
A.2 Near- and far-field intensity distributions
The output facet of the BAL is imaged onto a CMOS camera (Allied Vision Mako U-503B) using a 25×25$\times$25 × microscope objective (NA $=0.5$) and a lens with $175$ mm focal length (Thorlabs LA1399-B) in a telescope configuration. A beam splitter and a third lens (Thorlabs LA1050-B, $100$ mm focal length) are used to obtain the Fourier transform of the image plane for parallel far-field measurements with a second camera (Allied Vision Mako U-130B). The far-field angle is calculated from the magnification of the telescope and the focal length of the third lens. Neutral-density filters in the beam path and the exposure times of the cameras are adjusted during the measurement to avoid both low signal strength and saturation.
Figure 7a shows the measured near-field intensity distributions, which are almost uniformly flat well below threshold ($I_{th}=160$ mA). Just above threshold, the intensity distribution exhibits a strong dip in the center with two peaks near the edges of the output facet. As the pump current increases, these two peaks move further towards the edges, and the center of the intensity distribution slowly fills. Above approximately $850$ mA the near-field distribution develops a top hat profile with significant modulation (cf. Fig. 9). A similar evolution of the near-field intensity distributions as function of the pump current was observed in Ref. [64] and attributed to the interplay of index guiding and carrier-induced index changes. It should be noted that the near-field images are very symmetric around $x=0$ even though the individual transverse mode profiles show significant asymmetries (see Fig. 2c.).
The far-field intensity distributions shown in Fig. 7b are very broad below threshold and collapse to a single lobe at $0^{\circ}$ just above threshold. In this current regime, mainly the 1st- and 2nd-order modes lase, but the near-field distributions exhibit a pronounced dip in the center nonetheless since the fundamental mode features a dip in the center similar to the $q=2$ mode (cf. Fig. 2c). As the pump current increases, higher-order modes start to lase, and the power shifts away from the low order modes.
A.3 Time-domain measurements
Time-domain measurements were performed with a fast photodetector (Newport 1484-A-50, $15$ kHz to $22$ GHz) connected to a traveling-wave amplifier (Newport 1421, $8.5$ dB) and an oscilloscope (Tektronix DPO72304SX, $23$ GHz). Time traces were measured with $50$ GSample/s ($20$ ps sampling step) and $10^{6}$ points ($20~{}\mu$s long). RF-spectra $P_{RF}(f)$ are obtained by calculating the modulus squared of the Fourier transform $\mathcal{F}$ of the time traces $U(t)$ and smoothing with a moving average,
$P_{RF}(f)=\mathcal{M}\{|\mathcal{F}[U(t)]|^{2}\}\,,$ (1)
where $\mathcal{M}$ is the moving average over width $f_{\mathrm{avg}}$. A moving average width of $f_{\mathrm{avg}}=2$ MHz was used unless otherwise noted.
It should be noted that extremely narrow peaks are observed in many RF-spectra: these are artifacts due to synchronization issues between different analog-to-digital converters of the oscilloscope and appear even without input signal (see background signal in Fig. 2e). Hence they have no physical significance and can be ignored.
For measurements of the total laser emission, the entire beam was coupled into a graded-index multimode fiber (MMF, Thorlabs M116L02, $50~{}\mu$m diameter, NA$=0.20$, $2$ m long) using a 40×40$\times$40 × aspheric objective lens (Newport 5722-B-H, NA = $0.6$) and a fiber coupler (Thorlabs MBT613D/M). Feedback from reflections at the fiber facet were avoided by means of an optical isolator (Thorlabs IO-5-850-VLP). The multimode fiber was connected to the photodetector. It should be noted that coupling the beam of a multi-transverse mode laser into a single-mode fiber is very inefficient, and more importantly the higher-order transverse modes have a lower coupling efficiency compared to low-order modes, which results in an observational bias (see Appendix E).
A.4 Low-resolution optical spectra
@@
Figure 8 shows lasing spectra of the BAL measured with an optical spectrum analyzer (OSA, Anritsu MS9740A) with $30$ pm ($12.5$ GHz) resolution. The entire beam was coupled into the multimode fiber (see Appendix A.3). While the spectral resolution of the OSA is relatively low, it allows measurements over a broader spectral range than the heterodyning technique. Due to the low resolution, the spectra feature a single series of peaks with distance equal to the longitudinal-mode spacing (free spectral range, FSR) of $25.9$ GHz ($62.4$ pm), hiding the fact that the laser exhibits several transverse modes. As the pump current increases, the center of mass (COM) of the spectrum red-shifts (see Fig. 8b) due to Joule heating. The total shift is about $2$ nm over $800$ mA. The red-shift of the individual peaks in the spectrum can also be seen in Fig. 8a.
@@
A.5 Low-resolution spatio-spectral images
Figure 9: Spatio-spectral near-field image at 1000 mA measured with an imaging spectrometer. The spatial and spectral projections are indicated in red. The small white bar indicates the free spectral range (FSR).
Spatio-spectral images with low resolution were measured with an imaging spectrometer (Princeton Instruments SpectraPro HRS-500 with $1800$ g/mm grating, $30$ pm resolution) and a CCD camera (Allied Vision Mako U-503B). In order to measure the spectrally-resolved near-field intensity distributions, the output facet of the BAL was imaged onto the entrance slit of the spectrometer with 40×40$\times$40 × magnification. A dove prism (Thorlabs PS992M-B) was used to rotate the image of the BAL facet by $90^{\circ}$ so it matches the vertical orientation of the entrance slit.
The spatio-spectral image for $1000$ mA in Fig. 9 shows spatial patterns varying with the wavelength due to the presence of several transverse modes. However, the individual transverse modes cannot be resolved. The spectrally-integrated near-field intensity distribution has a top-hat profile with regular modulation which indicates the presence of high-order transverse modes (see Fig. 7 for more details). The spatially integrated spectrum shows only a single series of peaks with spacing equal to the FSR like the measurements with the OSA in Fig. 8, obfuscating the existence of several transverse modes.
@@ B.1 Experimental setup and measurement protocol
Figure 10: Experimental setup for spatially-resolved heterodyne measurements. a, Free-space part of the setup, where $x$ indicates the lateral position of the fiber. BAL: broad-area laser, Obj: 40×40$\times$40 × objective lens, OI: optical isolator, MMF: multimode fiber. b, Fiber-optic part of the setup. Single-mode fibers (multimode fibers) are indicated in yellow (orange), and coaxial cables in blue. TL: tunable laser (reference laser), eVOA: electronic variable optical attenuator, $\lambda$-meter: wavelength meter, PD: fast photodetector, RF-amp: RF-amplifier, OSA: optical spectrum analyzer, Scope: oscilloscope
The experimental setup is presented in Fig. 10. A 40×40$\times$40 × aspherical objective lens (Newport 5722-B-H, NA = $0.6$) and a plano-convex lens (Thorlabs LA1979-B, focal length $200$ mm) image the output facet of the BAL with 40×40$\times$40 × magnification in the plane where its emission is collected with a MMF (Thorlabs M116L02, see above). An optical isolator (Thorlabs IO-5-850-VLP) avoids parasitic feedback into the BAL from reflections at the fiber facet. The fiber is mounted on a motorized translation stage (Thorlabs NRT100/M with driver BSC201) and scans the transverse position on the output facet of the BAL with a step size of $20~{}\mu$m (141 positions). The corresponding step size on the BAL output facet, $0.5~{}\mu$m, is slightly smaller than the spatial resolution of $0.7~{}\mu$m estimated from the NA of the objective lens.
The signals from the BAL and the tunable reference laser (Toptica DL pro with Diode AR 860 nm) are superimposed via a multimode fiber-optic coupler (Thorlabs TM50R5F2B) on a fast photodetector and measured with an oscilloscope (see above). The optical powers of the reference laser and the BAL on the photodetector are adjusted to about $0.75$ mW and maximally $0.5$ mW, respectively, using a neutral density filter and a variable optical attenuator (VOA), though the power of the BAL can be smaller for low pump currents. The reference laser is set to $20^{\circ}$ with $143$ mA pump current and its wavelength adjusted to the desired value. It is left to stabilize for at least $20$ minutes, and we check that it is in a stable single-mode lasing regime. Its wavelength is monitored using a wavelength meter (High Finesse Angstrom WS6-600 IRVIS) connected via a $99$:$1$ fiber coupler (Thorlabs TW850R1A2).